OpenStudy (anonymous):

I WILL FAN AND MEDAL!!! Which of these ordered pairs is a solution to the equation y equals 2 x minus 1? Choose one answer. A. zero comma 1 B. 1 comma negative 3 C. 2 comma 3 D. 5 comma 3

OpenStudy (anonymous):

D :) \[y=2x-1 \] \[(5)=2(3) - 1\] \[5=6-1 \] \[5\]